Given the population 1,3,4,6 and 8. Suppose samples of size 3 are drawn from this population. How many different samples can be drawn from this population?



1
Expert's answer
2022-05-12T03:22:26-0400

Consider the samples without replacement (order of elements in samples does not matter). There are "C_5^3=\frac{5!}{3!2!}=10" different samples. Consider samples with replacement(order of elements in sample matters). There are "5^3=125" different samples. We used basic facts from combinatorics to compute a number of samples in both cases.

Answer: there are "10" different samples without replacement (order of elements in samples does not matter) and "125" different samples with replacement (order of elements in samples matters).